Arthur Andersen Enterprise Group/National Small Business United,Washington,
conducted a national survey of small-business owners to determine the challenges
for growth for their businesses. The top challenge, selected by 46% of the
smallbusiness owners, was the economy. A close second was finding qualified workers
(37%). Suppose 15% of the small-business owners selected both the economy and
finding qualified workers as challenges for growth. A small-business owner is
randomly selected.

(a) Given that the owner does not select the economy as a challenge for growth,
what is the probability that the owner believes that finding qualified workers is a
challenge for growth?
(b) What is the probability that the owner believes neither that the economy is a
challenge for growth nor that finding qualified workers is a challenge for growth?
1
Expert's answer
2019-07-25T14:00:13-0400

Let event A = small-business owner selected economy,

event B = small-business owner selected finding qualified workers, then

probabilities


"P(A)=46/100, P(B)=37/100,P(A\cap B)=15/100."

b) Using inclusion-exclusion principle


"P(A\cup B) = P(A)+P(B)-P(A\cap B)="

"46\/100+37\/100-15\/100 = 68\/100"

then  the probability that the owner believes neither that the economy is a

challenge for growth nor that finding qualified workers is a challenge for growth



"P(\bar{A}\cap \bar{B})= 1-P(A\cup B)=1 - 68/100 = 32/100."

a) Probability


"P(B\cap \bar{A})=P(B)-P(A\cap B)="

"37\/100 - 15\/100 = 22\/100."

Conditional probability that the owner believes that finding qualified workers is a

challenge for growth, given that the owner does not select the economy as a

challenge for growth is


"P(B|\bar{A})=P(B\cap\bar{A})/P(\bar{A})="

"(22\/100)\/(1-46\/100) = 22\/54 = 11\/27."

Answer: a) 11/27, b) 32/100.